# ANVIL: Anomaly-based Vulnerability Identification without Labelled Training Data

### v1

@@ VI Related Work
Anomaly Detection. Previous work has used anomalous behaviour as a signal for buggy code. Yun et al.  [30] detects API misuse by automatically modelling correct usage patterns from existing code. They use this model to detect anomalous uses, and hence bugs. Other work from Ahmadi et al.  [31] operates on singular codebases and clusters functionally similar yet inconsistent code. Their insight is that small inconsistencies between code that should have the same functionality can lead to bugs (e.g. missing checks). To our knowledge, the only previous attempt to leverage LLMs for anomaly-guided vulnerability detection used a simple string distance score between the LLM prediction and ground truth to determine if a line of source code is anomalous [32]. A string distance calculation lacks any contextual understanding of the strings being compared, and ANVIL improves upon this with a cross-entropy based anomaly measure. Further, their method of detecting anomalies relies heavily on pre-existing source code comments; our approach makes no assumptions on the contents of the software under examination. Finally, LLM-based anomaly detection in other non-code domains has also been explored, especially in log analysis and network intrusion detection [33].
@@
LLMs. Following the success of natural language generation with LLMs, work such as CodeBERT [1] and CodeT5 [2] have applied transformers toward code generation and understanding. CodeBERT is an encoder-only transformer that supports bidirectional natural language-programming language generation. However, it requires an additional decoder layer that cannot benefit from pre-training. To address this, Wang et al. proposed CodeT5, an encoder-decoder transformer. CodeT5 is pre-trained on tasks in which source code identifiers, and entire spans of code are masked for prediction. VulGen [13], extends previous work by training an LLM to identify and inject vulnerabilities into the codebase of programs for the purposes of synthetic data generation.
@@
ML-Assisted Dynamic Vulnerability Detection. Many existing fuzzers have leveraged machine learning to boost fuzzer performance. He et al.  [34] demonstrated a neural network’s ability to imitate the behaviour of a symbolic execution engine. They use a symbolic engine to generate input transactions to a corpus of smart contracts. They then train a neural network to predict these transactions. The network is then integrated into a fuzzer which they use to test unseen smart contracts. Recent work from Shi et al.  [35] uses a convolutional neural network to classify the behaviour of basic blocks. They then correlate these basic blocks with bytes in the program input. They use this mapping from input bytes to classification to guide an input-format-aware fuzzer. As for LLMs, a recent paper utilized ChatGPT to parse protocol specifications into a machine understandable grammar, which they then integrate into a protocol aware fuzzer [36].
@@
ML-Assisted Static Vulnerability Detection. To detect use-before-initialization bugs, Li et al. [37] augments a state-of-the-art static analysis tool with an LLM prompting technique. They utilize the LLM to extract variable initialization information when the static analysis tool is unable to handle certain code features (i.e. unsupported library calls). Conversely, Sun et al. [38] first utilizes an LLM to match candidate source code with potential vulnerabilities, then uses static analysis to verify its legitimacy. They apply their method to find vulnerabilities in smart contracts.
@@
ML-Based Vulnerability Detection. ML-based vulnerability detection largely focuses on the architecture and code features that will better lead a model to automatically detect vulnerabilities. An early approach from Russell et al. [11] attempts to find C/C++ vulnerabilities with a convolutional neural network. Further advances have experimented with graph neural networks [5] and long short-term memory models [4]. Unfortunately, on unseen bugs, these architecture level differences have not resulted in generalization to unseen, real-world vulnerabilities [39], and other work shows the same trend for prompt-guided LLM-based detectors [3]. Other work has focused on using different code features for training vulnerability detectors, such as code slices [10] and bug-triggering paths [7].
@@ VII Conclusion
To address the limitations of existing supervised-learning-based vulnerability detectors, we propose ANVIL, which leverages model mispredictions as guidance for software vulnerability detection. Our method capitalizes on the inherent capabilities of LLMs to generate normatively correct code, utilizing discrepancies in the model’s predictions to flag potential vulnerabilities. By employing code generation tasks across different LLMs, we have shown that this discrepancy in vulnerabilities is generalized to different model sizes and model architectures, and we also reveal that using compound statement contexts (MCS) outperforms simpler fix-sized contexts, providing a more relevant and precise framework for evaluating vulnerabilities. Furthermore, comparative evaluations against state-of-the-art supervised models LineVul and LineVD highlight ANVIL’s superior performance, achieving significantly better ROC-AUC and Top-5 accuracy metrics without the need for a labelled training dataset.

### v2

@@ VI Related Work
Anomaly Detection. Previous work has used anomalous behaviour as a signal for buggy code. Yun et al.  [yun_apisan_2016] detects API misuse by automatically modelling correct usage patterns from existing code. They use this model to detect anomalous uses, and hence bugs. Other work from Ahmadi et al.  [ahmadi_finding_2021] operates on singular codebases and clusters functionally similar yet inconsistent code. Their insight is that small inconsistencies between code that should have the same functionality can lead to bugs (e.g., missing checks). To our knowledge, the only previous attempt to leverage LLMs for anomaly-guided vulnerability detection used a simple string distance score between the LLM prediction and ground truth to determine if a line of source code is anomalous [ahmad_flag_2023]. A string distance calculation lacks any contextual understanding of the strings being compared, and ANVIL improves upon this with a cross-entropy-based anomaly measure. Further, their method of detecting anomalies relies heavily on pre-existing source code comments; our approach makes no assumptions on the contents of the software under examination. Finally, LLM-based anomaly detection in other non-code domains has also been explored, especially in log analysis and network intrusion detection [su_large_2024].
@@
LLMs. Following the success of natural language generation with LLMs, work such as CodeBERT [feng_codebert_2020] and CodeT5 [wang_codet5_2021] have applied transformers toward code generation and understanding. CodeBERT is an encoder-only transformer that supports bidirectional natural language-programming language generation. However, it requires an additional decoder layer that cannot benefit from pre-training. To address this, Wang et al. proposed CodeT5, an encoder-decoder transformer. CodeT5 is pre-trained on tasks in which source code identifiers, and entire spans of code are masked for prediction. VulGen [nong_vulgen_2023], extends previous work by training an LLM to identify and inject vulnerabilities into the codebase of programs for the purposes of synthetic data generation.
@@
ML-Assisted Dynamic Vulnerability Detection. Many existing fuzzers have leveraged machine learning to boost fuzzer performance. He et al.  [he_learning_2019] demonstrated a neural network’s ability to imitate the behaviour of a symbolic execution engine. They use a symbolic engine to generate input transactions to a corpus of smart contracts. They then train a neural network to predict these transactions. The network is then integrated into a fuzzer which they use to test unseen smart contracts. Recent work from Shi et al.  [shi_aifore_2023] uses a convolutional neural network to classify the behaviour of basic blocks. They then correlate these basic blocks with bytes in the program input. They use this mapping from input bytes to classification to guide an input-format-aware fuzzer. As for LLMs, a recent paper utilized ChatGPT to parse protocol specifications into a machine understandable grammar, which they then integrate into a protocol aware fuzzer [meng_large_2024].
@@
ML-Assisted Static Vulnerability Detection. To detect use-before-initialization bugs, Li et al. [li_enhancing_2024] augments a state-of-the-art static analysis tool with an LLM prompting technique. They utilize the LLM to extract variable initialization information when the static analysis tool is unable to handle certain code features (e.g., unsupported library calls). Conversely, Sun et al. [sun_gptscan_2024] first utilizes an LLM to match candidate source code with potential vulnerabilities, then uses static analysis to verify its legitimacy. They apply their method to find vulnerabilities in smart contracts.
@@
ML-Based Vulnerability Detection. ML-based vulnerability detection largely focuses on the architecture and code features that will better lead a model to automatically detect vulnerabilities. An early approach from Russell et al. [russell_automated_2018] attempts to find C/C++ vulnerabilities with a convolutional neural network. Further advances have experimented with graph neural networks [cheng_deepwukong_2021] and long short-term memory models [li_vuldeepecker_2018]. Unfortunately, on unseen bugs, these architecture level differences have not resulted in generalization to unseen, real-world vulnerabilities [chakraborty_deep_2022], and another work shows the same trend for prompt-guided LLM-based detectors [steenhoek_comprehensive_2024]. Other work has focused on using different code features for training vulnerability detectors, such as code slices [mirsky_vulchecker_2023] and bug-triggering paths [cheng_how_2022].
@@ VII Conclusion
To address the limitations of existing supervised-learning-based vulnerability detectors, we propose ANVIL, which leverages model mispredictions as guidance for software vulnerability detection. Our method capitalizes on the inherent capabilities of LLMs to generate normatively correct code, utilizing discrepancies between the model’s code predictions and the actual code, to flag potential vulnerabilities. By employing code generation tasks across different LLMs, we have shown that this discrepancy is generalizable to different model sizes and model architectures, and we also reveal that using Adaptive Context (AC) outperforms simpler fix-sized contexts, providing a more relevant and precise framework for evaluating vulnerabilities. Furthermore, comparative evaluations against state-of-the-art supervised models LineVul, LineVD and LLMAO highlight ANVIL’s superior performance, achieving significantly better vulnerability identification without the need for a labelled training dataset.
Appendix A Vulnerability Classification & Prioritization on PrimeVul using LLMAO Trained on Devign
[…]
LLMAO (6B) Devign 5.9% 14.0% 25.8% 43.7 0.34 51.3%
LLMAO (16B) Devign 4.4% 14.0% 22.5% 50.7 0.40 50.7%
Appendix B Distribution of CWEs in PrimeVul Validation and Test sets
CWE Name Line Count
1 CWE-787 Out-of-bounds Write 174
2 CWE-703 Improper Check or Handling of Exceptional Conditions 168
3 CWE-125 Out-of-bounds Read 101
4 CWE-416 Use After Free 88
5 CWE-476 NULL Pointer Dereference 77
6 CWE-369 Divide By Zero 56
7 CWE-400 Uncontrolled Resource Consumption 45
8 CWE-190 Integer Overflow or Wraparound 30
9 CWE-20 Improper Input Validation 29
10 CWE-119 Improper Restriction of Operations within the Bounds of a Memory Buffer 24
11 CWE-120 Buffer Copy without Checking Size of Input 23
12 CWE-354 Improper Validation of Integrity Check Value 18
13 CWE-617 Reachable Assertion 16
14 CWE-200 Exposure of Sensitive Information to an Unauthorized Actor 14
15 CWE-189 Numeric Errors 14
16 CWE-444 Inconsistent Interpretation of HTTP Requests 13
17 CWE-754 Improper Check for Unusual or Exceptional Conditions 13
18 CWE-362 Concurrent Execution using Shared Resource with Improper Synchronization 12
19 CWE-310 Cryptographic Issues 6
20 CWE-415 Double Free 6
21 CWE-59 Improper Link Resolution Before File Access 6
22 CWE-287 Improper Authentication 4
23 CWE-401 Missing Release of Memory after Effective Lifetime 3
24 CWE-843 Access of Resource Using Incompatible Type 3
25 CWE-191 Integer Underflow (Wrap or Wraparound) 2
26 CWE-276 Incorrect Default Permissions 2
27 CWE-399 Resource Management Errors 2
28 CWE-79 Improper Neutralization of Input During Web Page Generation 2
29 CWE-295 Improper Certificate Validation 2
30 CWE-668 Exposure of Resource to Wrong Sphere 1
31 CWE-704 Incorrect Type Conversion or Cast 1
32 CWE-665 Improper Initialization 1
33 CWE-269 Improper Privilege Management 1
34 CWE-345 Insufficient Verification of Data Authenticity 1
35 CWE-824 Access of Uninitialized Pointer 1
36 CWE-122 Heap-based Buffer Overflow 1
37 CWE-924 Improper Enforcement of Message Integrity During Transmission in… 1
38 CWE-908 Use of Uninitialized Resource 1
39 CWE-134 Use of Externally-Controlled Format String 1
40 CWE-284 Improper Access Control 1
Total 964